Show that the wavelength of electromagnetic radiation is equal to the de Broglie wavelength of its quantum (photon)

Grade:11

1 Answers

Vikas TU
3 years ago
The momentum of an electromagnetic wave of frequency V, wavelength λ is given by
p = hv /c = h / λ
or
λ = h/p
de broglie wavelength of proton, λ = h/p
h/p = h/ (hv/c) = c/v
Thus, wavelength of electromagnetic radiation is equal to the de-Broglie wavelength.